Circular No. 9000

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 U1 = P/2000 S7 (McMILLAN) S. Nakano, Sumoto, Japan, reports his identification of two nights of

previously unpublished observations (here designated P/2000 S7), obtained with the Lowell Observatory 0.59-m Schmidt telescope in the course of the LONEOS project (astrometry given below), with comet P/2008 U1 (cf. IAUC 8997).

2000 UT  α_{2000}  δ_{2000}  Mag.
Sept.23.20248  23^{h}09^{m}38.^{s}53  −9^{o}03′32′′.6  18.1
23.21798  23 09 37.84  −9 03 39.7
23.23342  23 09 37.45  −9 03 43.4
23.24886  23 09 36.96  −9 03 50.6
Oct. 2.18685  23 05 32.41  −9 50 41.0  18.3
2.20628  23 05 32.03  −9 50 44.1
2.22565  23 05 31.36  −9 50 51.9
2.24506  23 05 31.00  −9 50 57.4

These observations required a correction of ΔT = +1.19 days in 2000 from an orbit utilizing astrometry from 2008 Sept. 20–Nov. 21. Nakano provides the following orbital elements linking 103 observations, 2000–2008 (mean residual 0′′.66). The comet passed 0.18 AU from Jupiter in 2004 July 8.

Epoch = 2000 Dec. 2.0 TT

T = 2000 Dec. 8.9745 TT  ω = 217.^{o}3768
e = 0.350672  Ω = 152.9971 }2000.0
q = 2.693367 AU  i = 7.1433

a = 4.147928 AU n^{o} = 0.1166695 P = 8.45 years Epoch = 2008 May 14.0 TT

T = 2008 May 13.2573 TT  ω = 310.^{o}4836
e = 0.374407  Ω = 36.4178 }2000.0
q = 2.524827 AU  i = 4.4128

a = 4.035895 AU n^{o} = 0.1215610 P = 8.11 years

COMET C/2006 W3 (CHRISTENSEN) Visual total-magnitude estimates: Sept. 7.81 UT, 10.4 (K. Yoshimoto,

Yamaguchi, Japan, 25-cm reflector); Oct. 18.81, 10.5 (A. Diepvens, Balen, Belgium, 15-cm refr.); Nov. 16.74, 10.3 (B. Granslo, Fjellhamar, Norway, 20-cm refl.); 30.80, 10.0 (M. Meyer, Limburg, Germany, 15-cm refr.).

2008 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green